OpenStudy (anonymous):

I have 4 equations. I really just don't get the reciprocal. 1. x+9/3=8 2. x/4-6=-5 3. 5x/6+15=-10 4. x/2+3=13

zepdrix (zepdrix):

I can't quite tell with your problems since there are no brackets. Is this what we're dealing with?\[\Large\sf \frac{x+9}{3}=8\]Or this?\[\Large\sf x+\frac{9}{3}=8\]

OpenStudy (mathmale):

Important: Do you really mean "reciprocal," or do you mean "inverse function?"

OpenStudy (anonymous):

@zepdrix , the first one. @mathmale , I honestly don't know.

zepdrix (zepdrix):

\[\Large\sf \frac{x+9}{3}=8\]So we have a fraction... But you can think of the denominator as division. So we're dividing by 3. Inverse of division is multiplication. Let's multiply both sides by 3,\[\Large\sf 3\cdot\frac{x+9}{3}=8\cdot 3\]

zepdrix (zepdrix):

On the left side, since we're dividing and multiplying by 3, these operations undo one another. Or you can think of them as "cancelling out" if that helps.\[\Large\sf \cancel3\cdot\frac{x+9}{\cancel3}=8\cdot 3\]8*3 simplifies to 24. So we're left with,\[\Large\sf x+9=24\]Then it's just a simple subtraction step to solve for x. Any of those steps really confusing?

OpenStudy (anonymous):

No. it made sense actually. How would I do number 3 though since the top number has both a number and a variable?

zepdrix (zepdrix):

This is number 3?\[\Large\sf \frac{5x}{6}+15=-10\]

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

Fractions aren't very fun to deal with. Let's get rid of the fraction on x before going further. We'll multiply both sides by the value in the denominator. So we'll multiply both sides by 6,\[\Large\sf 6\left(\frac{5x}{6}+15\right)=-10\cdot 6\]The reason I put brackets around the left side is so you remember to multiply the 6 by the 15 as well. You can't just get rid of the 6 under the x, you have to give a 6 to `each term` on the left side of the equation.

zepdrix (zepdrix):

Distributing the 6 to each term gives us,\[\Large\sf 6\cdot\frac{5x}{6}+15\cdot6=-10\cdot 6\]The 6/6 cancel out as in the last problem, and then let's multiply out these other values,\[\Large\sf \cancel6\cdot\frac{5x}{\cancel6}+15\cdot6=-10\cdot 6\]giving us,\[\Large\sf 5x+90=-60\]

zepdrix (zepdrix):

Then it's just a few more steps to solve for x. Anything too confusing? Understand how to finish it from there? :o

OpenStudy (anonymous):

That made a lot of sense! Thank you.

zepdrix (zepdrix):

cool \c:/

OpenStudy (triciaal):

the above is correct. it is always easier to work with variable on one side and numbers on the other when you can. using the #3 above 5x/6 + 15 = -10 5x/6 = -25 multiply each side by 6/5 (to make the coefficient of the variable 1) x = ?

OpenStudy (triciaal):

reciprocal use "upside down " of whatever is in front of the variable the product of a number and its reciprocal is 1 5/6 * 6/5 = 1